OpenStudy (anonymous):

which is an equation for the line with slope 1/3 through (-2,1) ? a.y=1/3x+1 b.y=1/3x+5/3 c.y=1/3x-5/3 d.y=1/3x+1/3

OpenStudy (jackellyn):

(y-y1)=1/3(x-x1) y-(1)=1/3(x-(-2)) y-1 = 1/3(x+2) y-1=1/3x + 1/3(2) y-1 = 1/3x + 2/3 y= 1/3x + 2/3 +1 combine like terms and that is your answer @jkev1

OpenStudy (jackellyn):

what do you get?

OpenStudy (anonymous):

.y=1/3x-5/3 @jackellyn ?